# secReferences in the Online Appendix

### v2

@@ Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ objectives and how their strategies evolve over time. Using data from 2,500 campaigns, we find that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date. Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer boycotts and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
@@ 1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels. Through tactics of “private politics,” such as boycotts and shaming campaigns, NGOs have effectively pressured companies to reform their practices (Baron, 2001; Egorov and Harstad, 2017). Many studies examine how campaigns by both large and small NGOs impact targeted firms, often showing enduring effects (e.g., Spar and La Mure, 2003; Yaziji and Doh, 2009). For instance, just weeks before McDonald’s 2006 Annual General Meeting (AGM), Greenpeace launched a campaign branding the company ‘McAmazon‘ for alleged links to Amazon deforestation, prompting McDonald’s to adopt a new sustainable sourcing policy within days (Goldberg and Droste Yagan, 2007). Similar campaigns have targeted companies across various sectors, including Nike, Shell, Monsanto, and BNP Paribas. Despite the significance of these campaigns in shaping NGO strategies and influencing corporate behavior, we know relatively little about NGOs’ objectives, trade-offs, and how these evolve over the course of an NGO’s life.
Researchers addressing these questions face several challenges, particularly due to the scarcity of data on NGOs. For instance, donation data for non-political nonprofits is difficult to access because of donor privacy concerns. [^1] Data on donations to political parties and charities are more readily available due to disclosure rules and surveys (e.g., Cagé and Rouzet, 2015; Perez-Truglia and Cruces, 2017; Bouton et al., 2022; Guillot and Cagé, 2023). This paper differs from this literature in that it does not focus on the connection between NGO and politics, but rather on the incentives that NGOs face when deciding to campaign against a firm. NGOs are also subject to much less stringent transparency requirements than other firms such as publicly listed corporations, which complicates the tracking of their fund usage. Moreover, the absence of a quantitative framework to systematically analyze NGO objectives presents another significant challenge, as NGOs’ diverse missions cannot easily be reduced to profit maximization (Maxwell, 2012; Fioretti, 2022). To obviate to this problem, the literature emphasizes extreme behaviors of NGOs, with some contributing to the greater good (e.g., Daubanes and Rochet, 2019), while others are captured by corporate interests (e.g., Bertrand et al., 2021, 2023).
The starting point of our analysis is that NGO campaigns are strategically timed. Drawing on campaign data from a wide range of NGOs—spanning smaller entities like the ‘Physicians Committee for Responsible Medicine‘ and the ‘European Center for Digital Rights‘ to major players like ‘PETA‘ and the ‘Sierra Club‘—we identify two key factors that influence an NGO’s ability to capture stakeholders’ attention. First, media coverage amplifies NGO campaigns, making efforts like shaming and boycotting more effective. [^2] Consistent with this, Hawn (2021) finds that media coverage of a firm’s irresponsible behavior can significantly amplify negative public perceptions (Shea and Hawn, 2019), potentially resulting in the failure of planned merger. Second, NGOs with stronger reputations are better positioned to influence target firms, as they can leverage broader support from stakeholders such as workers, consumers, shareholders, suppliers, and the public. We demonstrate that these two factors create a trade-off regarding the strategic decision of campaign timing. NGOs must choose between maximizing influence on firms and enhancing their reputation.
Employing over ten years of campaign data against the largest US corporations, we find that NGOs can improve their reputation by launching campaigns near the target firm’s AGM. The AGM attracts the highest media coverage, as media outlets engage with managers and key shareholders to discuss the firm’s current achievements and future outlook (Dyck et al., 2008). This heightened media attention motivates NGOs to campaign during this period, increasing their visibility and positively influencing the perception held by the public and the target firm regarding the NGO’s stance on the issue, as well as the NGO’s commitment to driving change, thereby accumulating reputation (Bar-Isaac et al., 2008; Gent et al., 2015; Mitchell and Stroup, 2017). Since our estimation routine compares the timing of the campaigns of NGOs campaigning against multiple firms in a year, we can account for any confounding factors varying within an NGO (and target firms) across years through fixed effects, such as their size.
A late campaign is unlikely to influence shareholder votes at the AGM, as most shareholders have already voted by proxy (Bebchuk et al., 2020). Nevertheless, we find that the heightened visibility of campaigns timed with the AGM increases the likelihood of (i) a shareholder proposal on a related topic the following year and (ii) a boycott campaign in the subsequent months, compared to campaigns launched at other times. Campaigns initiated one to two months before the AGM are more likely to lead to successful votes on related proposals at the current AGM, revealing a new trade-off between influence and exposure.
We argue that NGOs navigate this tradeoff differently depending on their reputation. We first show that NGOs are more likely to campaign on AGM dates early in their history but gradually shift to launching campaigns two or three months prior the AGM to impact votes directly. Our analysis rules out learning this aspect of proxy voting as an explanation for our finding, as we observe the same trend for NGOs whose first campaign occurred just after an AGM, meaning they already knew they could not influence the vote by campaigning that late. We also find no evidence that NGOs select different target firms as they gain more campaigning experience.
Second, we demonstrate that the success of a shareholder proposal at a current AGM is linked to how many times the NGO backing the proposal has campaigned on AGM dates in the past. We interpret this as a reduced-form proxy for the reputation built by the NGO for fighting over issues. As we document diminishing returns to targeting AGM dates, once an NGO has accumulated sufficient reputation, it is more effective at influencing shareholder votes through early campaigns. Conversely, we find that accumulating experience at campaigning outside AGM dates does not help NGOs at swaying shareholder proposal votes.
Empirically, our approach leverages exogenous variation in the timing of AGM meetings, which in the U.S. we show to be relatively stable within firms over time due to regulations by the Securities and Exchange Commission (SEC) designed to give shareholders adequate time to review proposals before voting them. Across analyses, our approach also allows us to exclude time-varying confounding factors at the firm or industry, and campaign topic levels. In particular, while current events and news may influence an NGO’s decision to launch a campaign, they cannot influence the AGM calendar. As a result, objectives are identified by the tradeoff between targeting a firm’s AGM date to increase their exposure and impacting the firm’s decisions by campaigning early in support of the news or event.
Although we do not structurally estimate stakeholders’ beliefs about an NGO’s type (reputation)—an approach unfeasible with campaign data alone—our analysis indicates that the success of an NGO campaign in supporting a shareholder proposal at an upcoming AGM relies on the NGO’s past activism, highlighting the importance of reputation as a signaling mechanism (Bar-Isaac et al., 2008).
We interpret our results through the lens of firm life cycle models, where consumers initially lack information about a firm’s product quality, spurring firms to invest in reputation, especially when firms are young (e.g., Wernerfelt, 1988; Cabral, 2000, 2009). In this context, waiting for the target firm’s AGM is a costly investment for NGOs, as they must weigh the immediate recognition gained from protesting as soon as discovering a project they oppose against the benefits of waiting to protest when a firm is under substantial media scrutiny at a time that is salient for shareholders and decision makers.
Over time, NGOs may recognize that their campaigns garner significant support from key stakeholders like consumers and shareholders, reflecting a better reputation. Once established, they can leverage this reputation to directly impact firms’ decisions, engage in negotiations with firms, or spend their image to directly support shareholder proposals.
Taken together, our results describe a novel trade-off between exposure and (immediate) impact as a key driver of the societal consequences of NGO activism. Unlike our paper, these activists are often assumed to pursue social objectives, which may stem from prosocial behavior as described by Bénabou and Tirole (2006) or from moral preferences, according to Baron (2010). Their level of commitment can also depend on how closely they align with the cause, as noted by Besley and Ghatak (2005). Baron (2010) further illustrates that NGOs can emerge when socially motivated individuals coordinate efforts, drawing on frameworks outlined by Dixit (2004) and Tabellini (2008). When firms become targets of activist campaigns, firms may engage in “self-regulation” to preempt opposition and reduce potential harm. As firms’ capture of regulators has become an unavoidable feature of our economies, NGO activism surges to rebalance public and private interests (Glaeser and Shleifer, 2003; Egorov and Harstad, 2017; Daubanes and Rochet, 2019). [^3] A related literature highlights several instances in which NGOs can become subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021). For example, firms may strategically exploit their relationships with NGOs—through donations or partnerships—to further political agendas (Bertrand et al., 2020) or to regain consumer and investor trust after reputational damages (Akey et al., 2021).
@@
This paper also contributes to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; George et al., 2023; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear, even when CSR investments increase market value (e.g., Hawn and Ioannou, 2016; Hartzmark and Sussman, 2019; Hasan et al., 2022). [^4] A recent literature also suggests that government initiatives might even hinder private ones (Akey et al., 2024). The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016; Gine et al., 2017)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010; Elfenbein and McManus, 2010). However, other agents may play a pivotal role. Among them, NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Fioretti et al., 2023; Chuah et al., 2024; Koenig and Poncet, 2019).
Therefore, this short paper sheds new light on NGOs as coordinating various stakeholders such as shareholders and consumers, empirically highlighting a set of tradeoffs that arise from stakeholder involvement (Porter and Kramer, 2011; Kaplan, 2020). We provide new insights into the incentives NGOs face and how they can help align a firm’s action with the interests of its stakeholders (McGahan and Pongeluppe, 2023). In this direction, the results also suggest that other stakeholders can gain privately (i.e., reputation) by engaging with a firm, bringing empirical evidence to theories suggesting that stakeholders would not contribute resources to an organization without the potential to claim at least some of the value created (cf., Barney, 2018; McGahan, 2023).
Finally, our findings contribute to a growing body of literature that exploits the exogenous occurrence of NGO campaigns to asses changes in the performances and prosocial outcomes of targeted firms (e.g., Koenig and Poncet, 2022; Mazet-Sonilhac and Mésonnier, 2023; Conway and Boxell, 2024). Our study turns the table by focusing on the incentives that non-profit firms face when campaigning against for-profit firms.
@@
The paper is organized as follows. Section 2 outlines the data construction and provides summary statistics. The empirical approach is described in Section 3. Sections 4 and 5 examine media exposure and influence at the current AGM as potential drivers of AGM actions. Section 6 explores the role of NGO reputation in supporting related shareholder proposals. Section 7 discusses the results and Section 8 concludes.
2 Data and Summary Statistics
@@
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^5] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common topics are Environment/Pollution, Human & Labor Rights, Animal Rights, and Health.
@@
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the largest (by market capitalization) 3,000 US-listed firms’ Annual General Meetings (AGMs). For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
@@
We select environmental and social proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
@@
The following three sections examine the payoffs accruing to an NGO in connection with the timing of the campaigns it launched, as well as the reactions to the campaigns of central stakeholders such as consumers and shareholders. We present the key takeaways from each analysis at the end of each section. Next, we discuss the empirical approach in detail.
3 Empirical Strategy
Our empirical approach investigates the time gap between a firm’s AGM and the actions taken by NGOs or other stakeholders, while controlling for time-varying confounding factors at the firm, industry, and campaign topic levels. Importantly, our identification strategy relies on the independence of the corporate calendar from events that may motivate NGOs to launch campaigns. The timing of AGMs is largely exogenous, as SEC regulations mandate a stable schedule across years. Specifically, Rule 14a-8 ensures shareholders have ample time to review proposals before voting, requiring that shareholder proposals be filed at least 120 days before the proxy statement is released, based on the previous year’s AGM timing. This stabilizes AGM dates across consecutive years (Glac, 2014), as shown in Appendix Figure B2, which plots AGM months across adjacent years. Most variation outside the diagonal line—indicating significant fluctuations in AGM months—is attributed to the pandemic, as the SEC allowed firms to adjust AGM dates in 2020 to prevent COVID-19 spread and facilitate shareholder attendance (SEC, 2020). Our results remain robust even when excluding data from 2020 and 2021.
While current events may influence an NGO’s decision to launch a campaign against a firm, they are unlikely to affect the timing of AGMs. These external factors may shift an NGO’s objectives by altering incentives to target a firm. Our analysis leverages these shifts to identify a key trade-off for NGOs: whether to influence a firm’s strategies by launching campaigns early or to gain greater visibility by timing campaigns around the firm’s AGM.
Our results highlight NGO experience as a crucial factor in this trade-off. Before interpreting our findings, we present results in Appendix Table A1 suggesting that NGOs do not select targets based on their experience, namely the number of campaigns they launched in the past. Panel (c) reveals that both small and large NGOs—in terms of their number of campaigns in the sample—target firms with similar market capitalization and total assets. Targets also have similar numbers of yearly news mentions, environmental and social scores, and suffered a similar number of environmental and social campaigns. Panel (d) examines target selection by regressing each one of these firm-level variables against the number of past campaigns launched by the campaigning NGO, excluding (Column 1) or including (Column 2) fixed effects at the firm-by-year, NGO-by-year, and topic-by-year levels. This correlation analysis shows that NGOs’ campaign targets do not change as NGOs gain experience, although they are more likely to focus on environmental issues rather than social ones in later years due to the increasing prominence of environmental challenges over social ones.

### v3

@@ Abstract
We analyze the timing of NGO campaigns to shed light on NGOs’ strategies and how they evolve over time. Data from 2,500 campaigns show that NGOs are six times more likely to launch campaigns on their target’s Annual General Meeting (AGM) date. Although this strategy increases media exposure and stakeholder scrutiny, resulting in consumer backlash and related shareholder proposals at the following AGM of the targeted firm, it has no impact on current AGM votes. As NGOs build reputational capital, they adjust their timing to influence AGM votes, revealing the trade-offs they face to drive corporate change.
JEL classifications: L21, L31
Keywords: Non-Governmental Organizations, campaigns, social and environmental responsibility, ESG, firm dynamics, non-profit objectives
@@ 1 Introduction
Over recent decades, non-governmental organizations (NGOs) have emerged as influential social actors, shaping regulations and corporate strategies outside traditional political channels (Baron, 2001; Egorov and Harstad, 2017). However, their actions are often difficult to observe, and evidence on the incentives they face and how these shape their strategic decisions remains limited (Yaziji and Doh, 2009). Thus, NGOs are frequently perceived as passive actors, either selecting the most favorable action from a common good perspective (Daubanes and Rochet, 2019) or being dysfunctional and susceptible to corporate capture (Bertrand et al., 2020). This paper introduces a novel approach to analyzing how NGOs make decisions with implications for corporations and their stakeholders.
Consider an NGO with information about corporate wrongdoing that faces a strategic choice: launch an immediate campaign to foster dialogue or wait for the optimal moment when the firm is most sensitive to external pressure. While early action encourages engagement and potential change, strategic timing can amplify impact by signaling strong opposition among stakeholders.
A time when firms are particularly sensitive to external pressure is during their Annual General Meetings (AGMs), recurring shareholder meetings where strategic decisions are revealed and executives are held accountable. We show that NGOs are four to six times more likely to campaign on a target’s AGM date than at other times, a pattern driven by the heightened media coverage AGMs attract. This visibility helps NGOs reach stakeholders and the broader public while imposing reputational costs on firms. However, while AGMs provide a platform for NGOs to maximize exposure, their direct influence is limited—most shareholders vote before the meeting via proxy, reducing the immediate impact on voting outcomes (Bebchuk et al., 2020). Empirically, we leverage the exogenous timing of AGMs, which remains stable within firms due to SEC regulations, allowing us to focus on NGOs’ actions while holding firms’ strategic choices constant.
Our findings reveal the “life cycle” of NGO strategies (Wernerfelt, 1988; Cabral, 2000). Early-stage NGOs prioritize visibility by campaigning on AGM dates, despite their limited direct influence. We show that AGM-timed campaigns prompt consumer backlash and increase the likelihood of related shareholder proposals the following year. However, while these campaigns generate exposure, they do not directly drive change, as, for instance, the target’s environmental and social scores remain unchanged. Over time, NGOs targeting AGMs start campaigning before AGMs to maximize direct influence on corporate decisions. Early campaigns increase the likelihood of related shareholder proposals succeeding, ultimately driving measurable improvements in firms’ environmental and social performance.
Our interpretation is that AGM exposure can enhance NGO reputation. Supporting this view, we find that the success of a shareholder proposal on a given cause increases when an NGO campaigns on the same issue and year, with past AGM-date campaigns by the NGO amplifying this effect. However, the returns to AGM-date campaigns diminish as their frequency rises. In contrast, past NGO campaigns on non-AGM dates have no impact.
We view AGM-date campaigns as a way for NGOs to signal their type to stakeholders and build their “reputation,” as evidenced by the diminishing returns of AGM-targeting over time, which is consistent with reputation models (Bar-Isaac and Tadelis, 2008). [^1] This practice also generates additional negative news to a target firm only for an NGO’s first few campaigns. Later campaigns inflict no additional damage. Thus, the NGO may be willing to incur a cost to build its “reputation-for-toughness,” which it can then leverage in future campaigns (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). Hence, more mature NGOs shift to campaign before AGMs. Furthermore, NGO reputation is likely multidimensional, shaped both by a sustained track record of activism and by a demonstrated commitment to a cause. As such, NGOs may transition to early campaigning not only due to diminishing returns on visibility but also to establish credibility through constructive engagement with key stakeholders (Baur and Palazzo, 2011; Mitchell and Stroup, 2017).
By comparing NGO campaign outcomes across firms and issues within years, our analysis accounts for factors that vary across years within NGOs, firms, and issues, such as organizational competence, learning, moral authority, resources, and the salience of specific issues. NGOs also engage firms through partnerships and other channels, but since such activities are voluntarily disclosed—and thus subject to selection biases (Burger and Owens, 2010)—we rather focus on campaign timing, which must be public to be effective and, thus, observable to researchers.
Our findings open new avenues for research, as future work may empirically explore how NGOs choose between campaigning and other activist tactics or whether these different approaches act as complements or substitutes. The exogenous timing of AGMs provides a useful identification strategy for such analyses, as it holds the target firm’s strategic choices constant, allowing researchers to isolate NGOs’ choices and their incentives.
Our results reveal a novel tradeoff between exposure and direct impact as a key driver of the societal consequences of NGO activism (Briscoe and Gupta, 2016). Unlike our framework, much of the existing literature assumes that activists primarily pursue social objectives, either stemming from prosocial behavior (Besley and Ghatak, 2005; Bénabou and Tirole, 2006) or moral preferences (Baron, 2010), without discussing their incentives.
At the same time, several studies highlight instances in which NGOs themselves are subject to firm capture (e.g., Grossman and Helpman, 1994, 2001; Bertrand et al., 2014; Drutman, 2015; Bertrand et al., 2021). Firms may also strategically leverage their relationships with NGOs—through donations or partnerships—to advance political agendas (Bertrand et al., 2020), or to restore consumer and investor trust following reputational damage (Akey et al., 2021).
Our main contribution in this literature is to show that NGOs are not passive actors. Rather, their tactics evolve over time, shaping firms and their stakeholders’ responses in different ways depending on the strategies they employ (e.g., Hendricks and Sorensen, 2009; Cabral and Hortacsu, 2010). Thus, understanding NGOs’ incentives is essential to learn why firms internalize their externalities.
@@
We also contribute to the growing body of literature on firms’ objectives and their societal impact (Flammer, 2018; Fioretti, 2022; Fangwa et al., 2024). Firms can help address significant global challenges such as pollution and inequality by internalizing their externalities. However, the mechanisms by which firms do so without explicit public policies, like Pigouvian taxes, remain unclear (e.g., Hawn and Ioannou, 2016). The literature emphasizes that shareholders can influence firms through either voicing concerns—especially through voting at AGMs (Cuñat et al., 2016)—or selling shares (Broccardo et al., 2022; Bonnefon et al., 2022; Hart et al., 2022; Saint-Jean, 2024), whereas consumers exert influence through their willingness to pay (e.g., Gneezy et al., 2010). However, other agents may play a pivotal role. Among them, we highlight that NGOs can serve as a coordination mechanism among stakeholders with potentially unaligned preferences and empower them to demand change (Koenig and Poncet, 2019; Fioretti et al., 2023; Chuah et al., 2024; Mazet-Sonilhac and Mésonnier, 2023; McGahan, 2023).
@@
The paper is organized as follows. Section 2 outlines the data construction and provides summary statistics. The empirical approach is described in Section 3. Sections 4 and 5 examine media exposure and influence at the current AGM as potential drivers of AGM actions. Section 6 explores the role of NGO reputation in supporting related shareholder proposals. Section 7 discusses the results and concludes.
2 Data
@@
Our dataset integrates three sources. First, we use data from Sigwatch, a consulting firm tracking global NGO campaigns against corporations (Hatte and Koenig, 2020). The data includes the campaign’s public release date and cause (e.g., “GMOs in food”), the targeted company and its prominence (e.g., if mentioned in the headline), and the sentiment (varying from negative, -2, to positive, +2), which is generally negative. We focus on campaigns against US-listed firms with negative sentiment, where only one firm is mentioned, and it appears in the headline. [^2] Since our analysis links campaign timing with a firm’s AGM date, we exclude campaigns against multiple firms as their AGM dates may differ. The dataset spans 2,512 NGO campaigns launched by 731 unique NGOs and targeting 523 unique firms. The most active NGOs in our sample include well-known organizations such as the Sierra Club, PETA, and Greenpeace. The most targeted companies are Monsanto, McDonald’s, Facebook, Coca-Cola, Apple, and Exxon Mobil. The most common causes are “environment/pollution,” “human & labor rights,” “animal rights,” and “health.”
@@
Second, we use data from Institutional Shareholders Service (ISS), a key provider of corporate voting data, on shareholder proposals at the Annual General Meetings (AGMs) of the 3,000 largest US-listed firms by market capitalization. For those firms, ISS data report detailed information on shareholder proposals for these firms, including proposals withdrawn before the AGM.
@@
We select environmental and social (E&S) proposals (4,483 in total) from the complete set of proposals filed at US firms between 2010 and 2021 (837 firms in total). Around a third of these E&S proposals are successful: either withdrawn from the voting agenda before the AGM or passed after going through a shareholder vote. [^3] Evidence shows that shareholders agree to withdraw the proposal only if the company commits to its implementation (Mueller et al., 2023). We document two facts: the number of E&S shareholder proposals has increased by 40% between 2010 and 2021, and their success rate has increased from 35% in 2010 to almost 50% in 2021. Appendix Figure B1 illustrates these trends.
@@
The following three sections examine an NGO’s payoffs in connection with the timing of its campaigns, and the reactions of central stakeholders. We present the key takeaways from each analysis at the end of each section. Before that, we detail our empirical approach.
3 Exploiting AGM Dates
Our empirical approach examines the timing of NGO campaigns relative to AGMs, leveraging the exogeneity of AGM dates, which remain stable within firms across years due to SEC regulations. Rule 14a-8 mandates shareholder proposals be filed 120 days before the proxy statement, anchoring AGM dates across years (Glac, 2014). Appendix Figure B2 confirms this stability, with exceptions largely due to pandemic-related adjustments allowed by the SEC (2020) to facilitate shareholder attendance at AGMs. [^5] Results remain robust excluding the covid period (2020–2021). While external events may influence NGOs’ campaign decisions, they do not affect AGM timing. Thus, AGMs allow us to isolate NGOs’ strategic trade-off from firms’ responses.
Target selection. Our analysis focuses on campaign timing, abstracting from NGOs’ selection of a target, which we leave for future research. However, Appendix Table A1 shows that NGOs’ targets does not depend on prior campaign experience, a key variable of our analysis. Panel (c) finds no differences in market capitalization, assets, news mentions, ESG scores, or campaign history between targets of small and large NGOs. Panel (d) regresses firm characteristics on NGOs’ past campaigns, showing no correlation between target selection and NGO experience. However, we see that NGOs increasingly focus on environmental over social issues in later years, reflecting the rising importance of environmental concerns.